OpenStudy (anonymous):

The base of a rectangle prism is 6in by 9in. What is the lateral area of the prism

ganeshie8 (ganeshie8):

are you given height ?

ganeshie8 (ganeshie8):

Lateral area of prism = (perimeter of base)*(height)

OpenStudy (anonymous):

I am going to be so tired for my final in the morning...9in by 6in by 5in

OpenStudy (anonymous):

ok so I figured out the total area its 258 cm now how do I figure lateral from there

OpenStudy (anonymous):

I told you sleep was important... =)

OpenStudy (anonymous):

I do horrible on tests when I haven't slept ><

OpenStudy (anonymous):

I know I got 26 more questions and then bed I have to do this or I will fail

ganeshie8 (ganeshie8):

you're far from failing, with all the hard work you have been putting :)

ganeshie8 (ganeshie8):

remember this formula for any prism : Lateral area of prism = (perimeter of base)*(height)

OpenStudy (anonymous):

I am at a 70 percent and I don't want to fail so I got to figure out what I don't know how to do

ganeshie8 (ganeshie8):

9in by 6in by 5in perimeter of base = 2(9+6) = 2(15) = 30 height = 5 plug them in the formula

ganeshie8 (ganeshie8):

Lateral area of prism = (perimeter of base)*(height) = (30) * (5) = 150

OpenStudy (anonymous):

ok I got this one I can do this I know I can thank you so much for the help I would give you 30 medals if I could